OpenStudy (anonymous):

pleaseee helpp !!!!what is the period and amplitude of the function for y=3 sin 4x

OpenStudy (anonymous):

amplitude is the absolute value of the number sitting out front

OpenStudy (anonymous):

period of \[\sin(bx)\] is \[\huge \frac{2\pi}{b}\]

OpenStudy (irishboy123):

have you tried drawing it? start with y = sin(x), then draw y = 3 sin(x), then the harder part draw y = 3 sin(4x). by draw, a simple doodle will suffice, and i am happy to help, but doing that really fires the synapses!

OpenStudy (ybarrap):

$$ y=A\sin(\omega t)\\ $$ Where \(\omega\) is the angular frequency: \(\omega = {2\pi\over T}\), \(T\) is the period and \(A\) is the Amplitude. So $$ \omega = b\\ \implies {2\pi\over T}=b\\ \implies T={2\pi\over b}\text{,which is the period.}\\ $$ It should be clear from this what the value of \(A\) is to make the following equal to your equation: $$ y=A\sin \left ({2\pi \over T}\cdot x \right)={\color{red}{A}}\sin bx={\color{red}{3}}\sin bx $$

OpenStudy (precal):

y=asin(b(x-c))+d a is your amplitude but it is the absolute vaue of a so example y=-2sin(4x) the amplitude is 2 your period is (2pi)/b in this case b is 4 just sub (2pi)/4 which is pi/2 can you apply this to your problem?

OpenStudy (precal):

d is the midline and c is your phase shift

OpenStudy (irishboy123):

@pham96 post this agin here, or on the physics forum, if this has been unhelpful. cheers!